Hartwick is considering leasing a new equipment. The lease lasts for 5 years. The lease calls...

  1. Hartwick is considering leasing a new equipment. The lease lasts for 5 years. The lease calls for 5 payments of $11,300 per year with the first payment occurring immediately. The equipment would cost $44,000 to buy and would be straight-line depreciated to a zero salvage value over 5 years. The actual salvage value is negligible because of technological obsolescence. The firm can borrow at a rate of 6%. The corporate tax rate is 25%. What is the NPV of the lease relative to the purchase?

    -$3,218.66

    -$4,537.23

    $737.11

    $2,912.64

    $1,097.26
Answer #1

NPV of lease = 11,300(1-25%)*[1+PVAF(6%, 4 years)

= 8,475*4.46510

=- $37,841.77

Annual Depreciation = 44,000/5 = $8,800

The firm will get tax savings each year on depreciation in buy

NPV = 8,800*25%*PVAF(6%, 5 years) – 44,000

= 2,200*4.21236 – 44,000

= -$34,732.80

Hence, NPV of the lease relative to the purchase = 34,732.80 – 37,841.77

= -$3,109

i.e. -$3,218.66 (approx.)